The Hartnett Corporation manufactures baseball bats with Pudge Rodriguezs autograph stamped on them. Each bat sells for $53Eaton Tool Company has fixed costs of $421,400, sells its units for $92, and has variable costs of $49 per unit. a. Compute tInternational Data Systems information on revenue and costs is relevant only up to a sales volume of 114,000 units. After 11
Answer #1

Please do Upvote if you are served. Feel free to reach out in the comments

Cheers!!!

Note :As HOMEWORKLIB's policies, I have answered first four parts of the question. Please ask the remaining parts in a seperate question. I will be more than happy to help !!

Hartnett Corporation:

Answer:

Contibution Margin = $53 - $28 = $25

=> Breakeven point = $40,000/25 = 1600 bats

Profit of $19,000:

Contribution margin = $19000+$40,000 = $59,000

=> Sales = $59,000/25 = 2360 bats

Eaton Company:

Answer:

Contribution margin = $92 - $49 = $43

Breakeven units = $421,400/$43 = 9800 units

Fixed costs reduces to $330,000

=> Breakeven units =$330,000/(92-52) = 8250 units